# Magnetism and hidden quantum geometry in charge neutral twisted trilayer graphene

### v1

@@ Introduction
Moiré materials offer a promising platform for designing tunable simulators of strongly correlated systems. The mismatch of the lattice constants or misalignment of the atomically-thin stacked layers leads to superlattice periodicity, which, for specific twist angles, flattens the bands around the Fermi energy. The flat bands enhance the role of the electron-electron interactions, leading to the observation of insulating and superconducting phases [1, 2] for bilayer graphene twisted by the so-called magic angle [3]. Following this discovery, a significant number of moiré systems has been realized with different number of layers and sequences of twist angles, including, but not limited to, mirror-symmetric trilayers [4, 5, 6], helical trilayers [7], double bilayers [8] and pentalayers [9]. Some of these have been shown to host a number of correlated and topological states such as correlated insulators [1, 10, 11, 12, 13, 14, 15], quantum anomalous Hall effect [16], ferromagnetism [17] and a generalized Wigner crystal state [18].
@@
Mirror-symmetric magic angle twisted trilayer graphene (TTG) consists of a middle layer rotated by the magic angle relative to two aligned top and bottom layers [19]. The band structure of TTG features two flat bands (per spin and valley) near the Fermi level, alongside Dirac cones that intersect the gap at the K points [19, 20, 21, 22, 23, 24, 25, 26, 27]. These features are also observed in ARPES experiments [28]. The spectrum of TTG is highly tunable through several external parameters, including the electric field, the hexagonal boron nitride (hBN) alignment, and the twist angle. Compared to twisted bilayer graphene (TBG), TTG offers a highly stable and versatile platform for engineering electronic states. Notably, TTG demonstrates robust and reproducible superconductivity across a wide range of tuning parameters [4, 29, 5, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Consequently, TTG has been the focus of an extensive theoretical investigation, employing both continuum models [19, 21, 22, 23, 24, 27] and atomistic approaches, such as ab initio [20, 40] and tight-binding [40]. Interaction effects in TTG have been explored in Refs. [23, 41, 26], while theoretical descriptions of its superconducting phase are proposed in Refs. [42, 43, 40, 44]. Additionally, the phonon properties of TTG were studied in Ref. [45]. Despite these significant advances, a unified theory of TTG is still lacking. Many of its properties, including the magnetic and topological phase diagrams, remain elusive.
@@
In this work, we consider the atomistic tight-binding Hamiltonian combined with a Hubbard interaction term and apply the self-consistent Hartree-Fock method to determine the magnetic properties of the ground state. Our primary focus is on the transition from the paramagnetic to the antiferromagnetic (AF) phase, a phenomenon extensively studied in monolayer graphene systems [46, 47, 48]. It has been shown that the area in the parameter space of the AF phase increases as we increase the number of layers [49]. We analyze the magnetic phase diagrams as functions of interaction strength, mirror, and sublattice symmetry-breaking perturbations, which arise from applied electric fields and hBN encapsulation. Interestingly, the interplay between magnetic ordering and topological properties, such as the emergence of AF Chern insulators [50], offers a compelling framework to explore quantum geometry in these systems. We compute the quasiparticle multiband Chern numbers for the valence and conduction flat bands, revealing the distribution of multiband Berry curvature (mBC) across the moiré Brillouin zone (mBZ), and demonstrate its tunability in experimentally realistic scenarios. We note that the behavior of mBC in TTG differs from the TBG case, where the $\Gamma$ point is the source of Berry’s curvature [51]. In the trilayer instance, we can identify contributions both from the $\Gamma$ and $K$ points.
@@ Results
Figure 1: Structural and electronic properties of TTG. (a) Side, (b) top, and (c) 3D view of geometry. TTG consists of three layers of graphene, with the middle one twisted by $\theta$, with respect to the aligned top and bottom layers, which are relaxed out-of-plane. The system is encapsulated in aligned hBN, and a perpendicular electric field is applied. (d) mBZ with high symmetry points. (e) Band structure of pristine TTG. (f) Effect of non-zero electric field ($\Delta_{V}\neq 0$) hybridizing Dirac cones with the flat band. (g) The combined effect of electric field and hBN encapsulation ($\Delta_{V}\neq 0,\Delta_{\rm hBN}\neq 0$), opening extra gaps near the Fermi level. The gaps opened by applying an electric field and hBN are marked with blue arrows.
@@
Figure 1 (a-c) summarizes our system’s geometry. The structure consists of top and bottom layers in AA stacking, with the middle layer twisted by an angle of $\theta=1.55^{\circ}$. The TTG is encapsulated in aligned hBN, which is modeled as a staggered potential applied to the top and bottom layers: $\hat{H}_{\Delta_{\rm hBN}}=\Delta_{\textrm{hBN}}\sum_{i,\sigma}\alpha_{i}|l_{i}|c_{i,\sigma}^{\dagger}c_{i,\sigma}$ where $l_{i}={1,0,-1}$ for top, middle and bottom layers respectively and $\alpha_{i}={1,-1}$ when $i$ belongs to sublattice A and B respectively. This way there is a $2\Delta_{\textrm{hBN}}$ difference in onsite energy between A and B atoms of the top and bottom layers. Furthermore, a perpendicular electric field is introduced, modifying the potential difference between the top and bottom layers: $\hat{H}_{\Delta_{V}}=\frac{\Delta_{V}}{2}\sum_{i,\sigma}l_{i}c_{i,\sigma}^{\dagger}c_{i,\sigma}$. Therefore, the full Hamiltonian is given by
@@
Since we assume a commensurate angle, i.e., periodic crystal structure, we can define a finite moiré unit cell (UC) (Fig. 1(b)) and mBZ (Fig. 1(d)). Each moiré UC consits of $N_{\rm at}$, and each atom repeats in the UC creating $N_{\rm at}$ simple Bravais sublattices. We apply periodic bondary condition to obtain allowed wavevectors, associate a Bloch wavefunction for a given wavevector $\vec{k}$ with each atom in a moiré cell, and diagonalize resulting Hamiltonian matrix to obtain energy bands. Fig.1(e) shows the result for the interaction strength $U=0$. This procedure is described in greater detail in Ref [52]. To find the low energy state of this Hamiltonian we use a self consistent Hartree-Fock procedure, see details in Methods. The resulting ground state is paramagnetic, meaning each atom has an equal spin population (with precision of $10^{-6}$). We obtain the characteristic flat band at the Fermi level with two valence and two conduction bands for each spin component. In our calculation, we also reproduce the high velocity Dirac cones at moiré K points, with Dirac points approximately 15 meV below the Fermi level. We confirm that the wave functions of these Dirac cones are localized solely on the top and bottom layers, while the flat band’s wave functions are localized $50\%$ on the middle layer and $25\%$ on the top and bottom (see Fig. 6 in SM). Such localization can be explained in terms of a TTG system comprising of a monolayer graphene and TBG [21].
We now study the effect of vertical electric field and hBN encapsulation. We first introduce a non-zero electric field ($\Delta_{V}\neq 0$). Such mirror-symmetry breaking perturbation hybridizes the Dirac cones with the flat bands [23]. This hybridization leads to splitting between the valence and conduction bands in the flat band, opening a gap at $K$ points, shown in Fig. 1(f) by a blue arrow. However, the gap at the Fermi level remains closed, and, in addition, the gap between the flat band and the remote bands closes at the $\Gamma$ point.
In the next step, we turn on the interaction with encapsulating hBN in the form of a top and bottom layer staggered sublattice potential Fig. 1(g). This leads to the opening of the gap within the flat band. Additionally, similar to the gapped monolayer graphene case [53], a gap between the Dirac cones opens and their dispersion becomes parabolic. The remote bands peel off from the flat band at the $\Gamma$ point so that our system consists of a well defined flat band and a gap around the Fermi energy. We note that TTG with an applied electric field and hBN is, therefore, strikingly similar to the TBG/hBN structure, with an extra set of remote bands at $K$ points originating from the gapped Dirac cones. More details of these mechanisms are available in SM.
@@
We now show the Hartree-Fock quasiparticle bands for spin up and down obtained by self-consistently diagonalizing Hubbard Hamiltonian, Eq. 4. For sufficiently small $U<U_{c}$, we do not observe any strong renormalization of the bands close to the Fermi level, as shown in Fig. 2(a). The state is paramagnetic, meaning each atom has an equal spin population. On the other hand, when $U>U_{c}$, a transition to an AF ground state occurs. A clear Mott gap is opened in the flat band after the phase transition in Fig. 2(b) for $U>1.1t$. There is no gap opening for the Dirac cones, therefore, no gap is generated at the Fermi level. The AF state has the most significant spin polarization in the middle layer, which is the layer with the highest contribution to the flat band. Bilayer graphene has smaller $U_{c}$ than a monolayer, therefore the effect of $U$ is stronger in the middle of TTG. The total spin density on A and B sublattices of graphene layer building TTG is imbalanced, producing a microscopic AF state. Such AF state is analogous to an AF state in other n-layer graphene systems [49, 54].
The Hubbard parameter $U$ for which the para- to anti-ferro- magnetic transition occurs is lower than in the mono- and bi-layer graphene systems (Fig. 3). For example, it has been shown that in the monolayer graphene (MLG) in the Hartree-Fock approximation $U_{c}\approx 2t$ [49]. We reproduced this result in our model, with small differences caused by the long-range TB hopping necessary to model TTG correctly. Equivalent calculations for bilayer graphene (BLG) yield $U^{\rm BLG}_{c}<U^{\rm MLG}_{c}$. This stems from the fact that since the dispersion in BLG is parabolic, the density of states is increased, and the ratio of interactions to kinetic energy increases as well. It is not surprising then that TBG has a smaller critical transition parameter ($U=1.1t$), as shown in Fig. 3.
@@
Multiband Berry’s curvature tuning. Flat bands of twisted materials are known to realize a variety of topological phases. For example, at $\nu=-3$ filling in TBG interaction-induced Chern insulator is predicted [17]. The topological phases catalog is still to be established in the TTG studied here. For a charge-neutral system for all studied electric fields, hBN strengths and Hubbard parameters $U$, flat bands near the Fermi level are always intertwined. Therefore, single band Berry’s curvature and Chern numbers are not well defined. However, multiband analogs can be defined, in our case of 2+2 bands separated from other bands by well-defined energy gaps, valid for an ample region in the parameter space of $\Delta_{\textrm{hBN}},\Delta_{V}$ and $U$. In Figs. 5 (a) and (b), we show the corresponding Hartree-Fock band structures for two representative choices of parameters (more details in SM). Details of multiband Berry’s curvature calculations are shown in the Methods section.
First, we establish that the multiband Chern number $C_{1+2+3+4}$ corresponding to all four bands is 0. This is also the case for the two flat valence (VBs) and conduction bands (CBs) separately ($C_{1+2}=C_{3+4}=0$). However, we observe a non-zero peak of mBC around $K$ and $K^{\prime}$ points, see e.g., Fig 5(c), which cancel each other out. We establish that non-zero peaks are also present around the $\Gamma$ point, in Fig. 5(c) in the corners of rhomboidal moiré mBZ. In that sense, the geometry of wavefunctions is hidden since it produces $C=0$ despite having a rich profile in the mBZ.
@@ Discussion
Twisted trilayer graphene is a highly tunable moiré system. We predict a surprisingly small critical value of Hubbard parameter $U$ necessary to observe magnetic phase transition between para- and anti-ferromagnetic states. This AF state is microscopic, in which spins are anti-aligned on neighboring graphene A and B sites inside TTG’s AAA moiré centers. It is well-known that the AF state is not observed in pristine graphene because interactions are too weak. On the other hand, in TTG, this condition is weakened, although we note that such a state is destabilized by mirror and sublattice symmetry breaking perturbations. Effectively, we also expect stronger screening in our setup compared to that of suspended graphene, and further studies are necessary to establish optimal device design to reach the AF state in TTG. Such studies, including, e.g., quantum fluctuations, more realistic extended Hubbard models with long-range interaction, and better account for the dielectric environment and electron-electron screening models are the subject of our ongoing work.
@@
Although topologically trivial at charge neutrality, TTG hosts nonzero peaks of multiband Berry’s curvature. This curvature originates from the geometry of wave functions in the flat band and can be tuned using electric fields. This can be useful in understanding low electron/hole doped phases, especially those in which carriers unequally populate valleys, i.e., valley polarized states. It would also be interesting to establish if AF and topological phases can co-exist for realistic interactions in other multilayer systems, e.g., pentalayers with hBN-induced moiré potentials, in which fractional quantum anomalous Hall states have been recently observed [55].

### v2

@@ Introduction
Moiré materials offer a promising platform for designing tunable simulators of strongly correlated systems. The mismatch of the lattice constants or misalignment of the atomically-thin stacked layers leads to superlattice periodicity, which, for specific twist angles, flattens the bands around the Fermi energy. The flat bands enhance the role of the electron-electron interactions, leading to the observation of insulating and superconducting phases [Cao_Jarillo-Herrero_2018a, Cao_Jarillo-Herrero_2018b] for bilayer graphene twisted by the so-called magic angle [Bistritzer_MacDonald_2011a]. Following this discovery, a significant number of moiré systems has been realized with different number of layers and sequences of twist angles, including, but not limited to, mirror-symmetric trilayers [Park_PJH_2021, Hao_Kim_2021, Cao_PJH_2021], helical trilayers [Devakul_Fu_2023], double bilayers [Su_Folk_2023] and pentalayers [Han_Ju_2024]. Some of these have been shown to host a number of correlated and topological states such as correlated insulators [Cao_Jarillo-Herrero_2018a, Tang_Fai_Mak_2020, Shen_Zhang_2020, Cao_PJH_2020, Liu_Kim_2020, Chen_Yankowitz_2021, Chen_Wang_2019], quantum anomalous Hall effect [Serlin_Young_2020], ferromagnetism [Sharpe_Goldhaber-Gordon_2019] and a generalized Wigner crystal state [Regan_Wang_2020].
@@
Mirror-symmetric magic angle twisted trilayer graphene (TTG) consists of a middle layer rotated by the magic angle relative to two aligned top and bottom layers [Khalaf_Vishwanath_2019]. The band structure of TTG features two flat bands (per spin and valley) near the Fermi level, alongside Dirac cones that intersect the gap at the K points [Khalaf_Vishwanath_2019, Carr_Kruchkov_2020, Calugaru_Bernevig_2021, Lei_MacDonald_2021, Xie_Lian_2021, Phong_Guinea_2021, Ledwith_Vishwanath_2021, Christos_Scheurer_2022, Yu_DasSarma_2023]. These features are also observed in ARPES experiments [Li_Liu_2022]. The spectrum of TTG is highly tunable through several external parameters, including the electric field, the hexagonal boron nitride (hBN) alignment, and the twist angle. Compared to twisted bilayer graphene (TBG), TTG offers a highly stable and versatile platform for engineering electronic states. Notably, TTG demonstrates robust and reproducible superconductivity across a wide range of tuning parameters [Park_PJH_2021, Cao_Jarillo-Herrero_2021, Hao_Kim_2021, Kim_Nadj-Perge_2022, Lin_Li_2022, Liu_Li_2022, Park_Jarillo-Herrero_2022, Siriviboon_Li_2022, Shen_Efetov_2023, Mukherjee_Deshmukh_2024, Batlle-Porro_Koppens_2024, Zhang_Li_2024, Zhou_Banerjee_2024]. Consequently, TTG has been the focus of an extensive theoretical investigation, employing both continuum models [Khalaf_Vishwanath_2019, Calugaru_Bernevig_2021, Lei_MacDonald_2021, Xie_Lian_2021, Phong_Guinea_2021, Yu_DasSarma_2023] and atomistic approaches, such as ab initio [Carr_Kruchkov_2020, Fischer_Klebl_2022] and tight-binding [Fischer_Klebl_2022]. Interaction effects in TTG have been explored in Refs. [Xie_Lian_2021, Qin_MacDonald_2021, Christos_Scheurer_2022], while theoretical descriptions of its superconducting phase are proposed in Refs. [Chou_DasSarma_2021, Lake_Senthil_2021, Fischer_Klebl_2022, Gonzalez_Stauber_2023]. Additionally, the phonon properties of TTG were studied in Ref. [Samajdar_Scheurer_2022]. Despite these significant advances, a unified theory of TTG is still lacking. Many of its properties, including the magnetic and topological phase diagrams, remain elusive.
@@
In this work, we consider the atomistic tight-binding Hamiltonian combined with a Hubbard interaction term and apply the self-consistent Hartree-Fock method to determine the magnetic properties of the ground state. Our primary focus is on the transition from the paramagnetic to the antiferromagnetic (AF) phase, a phenomenon extensively studied in monolayer graphene systems [Sorella_Tosatti_1992, Yazyev_2010, Guclu_Hawrylak_2014]. It has been shown that the area in the parameter space of the AF phase increases as we increase the number of layers [Vahedi_TramblydeLaissardiere_2021]. We analyze the magnetic phase diagrams as functions of interaction strength, mirror, and sublattice symmetry-breaking perturbations, which arise from applied electric fields and hBN encapsulation. Interestingly, the interplay between magnetic ordering and topological properties, such as the emergence of AF Chern insulators [Jiang_Wang_2018], offers a compelling framework to explore quantum geometry in these systems. We compute the quasiparticle multiband Chern numbers for the valence and conduction flat bands, revealing the distribution of multiband Berry curvature (mBC) across the moiré Brillouin zone (mBZ), and demonstrate its tunability in experimentally realistic scenarios. We note that the behavior of mBC in TTG differs from the TBG case, where the $\Gamma$ point is the source of Berry’s curvature [Zhang_Lu_2024]. In the trilayer instance, we can identify contributions both from the $\Gamma$ and $K$ points.
@@ Results
Figure 1: Structural and electronic properties of TTG. (a) Side, (b) top, and (c) 3D view of geometry. TTG consists of three layers of graphene, with the middle one twisted by $\theta$, with respect to the aligned top and bottom layers, which are relaxed out-of-plane. The system is encapsulated in aligned hBN, and a vertical electric field is applied. (d) mBZ with high symmetry points. (e) Band structure of pristine TTG. (f) Effect of non-zero electric field ($\Delta_{V}=120$ meV) hybridizing Dirac cones with the flat band. (g) The combined effect of electric field and hBN encapsulation ($\Delta_{V}=120$$\Delta_{\rm hBN}=25meV$), opening extra gaps near the Fermi level. The gaps opened by applying an electric field and hBN are marked with blue arrows.
@@
Figure 1 (a-c) summarizes our system’s geometry. The structure consists of top and bottom layers in AA stacking, with the middle layer twisted by an angle of $\theta=1.55^{\circ}$. The TTG is encapsulated in aligned hBN, which is modeled as a staggered potential applied to the top and bottom layers: $\hat{H}_{\Delta_{\rm hBN}}=\Delta_{\textrm{hBN}}\sum_{i,\sigma}\alpha_{i}|l_{i}|c_{i,\sigma}^{\dagger}c_{i,\sigma}$ where $l_{i}={1,0,-1}$ for top, middle and bottom layers respectively and $\alpha_{i}={1,-1}$ when $i$ belongs to sublattice A and B respectively. This way there is a $2\Delta_{\textrm{hBN}}$ difference in onsite energy between A and B atoms of the top and bottom layers. Furthermore, a vertical electric field is introduced, modifying the potential difference between the top and bottom layers: $\hat{H}_{\Delta_{V}}=\frac{\Delta_{V}}{2}\sum_{i,\sigma}l_{i}c_{i,\sigma}^{\dagger}c_{i,\sigma}$. Therefore, the full Hamiltonian is given by
@@
Since we assume a commensurate angle, i.e., periodic crystal structure, we can define a finite moiré unit cell (UC) (Fig. 1(b)) and mBZ (Fig. 1(d)). Each moiré UC consits of $N_{\rm at}$, and each atom repeats in the UC creating $N_{\rm at}$ simple Bravais sublattices. We apply periodic bondary condition to obtain allowed wavevectors, associate a Bloch wavefunction for a given wavevector $\vec{k}$ with each atom in a moiré cell, and diagonalize resulting Hamiltonian matrix to obtain energy bands. Fig.1(e) shows the result for the interaction strength $U=0$. This procedure is described in greater detail in Ref [Rodrigues_Hawrylak_2024]. To find the low energy state of this Hamiltonian we use a self consistent Hartree-Fock procedure, see details in Methods. The resulting ground state is paramagnetic, meaning each atom has an equal spin population (with precision of $10^{-6}$). We obtain the characteristic flat band at the Fermi level with two valence and two conduction bands for each spin component. In our calculation, we also reproduce the high velocity Dirac cones at moiré K points, with Dirac points approximately 15 meV below the Fermi level. We confirm that the wave functions of these Dirac cones are localized solely on the top and bottom layers, while the flat band’s wave functions are localized $50\%$ on the middle layer and $25\%$ on the top and bottom (see Fig. 6 in SM). Such localization can be explained in terms of a TTG system comprising of a monolayer graphene and TBG [Calugaru_Bernevig_2021].
We now study the effect of vertical electric field and hBN encapsulation. We first introduce a non-zero electric field ($\Delta_{V}=120$meV). Such mirror-symmetry breaking perturbation hybridizes the Dirac cones with the flat bands [Xie_Lian_2021]. This hybridization leads to splitting between the valence and conduction bands in the flat band, opening a gap at $K$ points, shown in Fig. 1(f) by a blue arrow. However, the gap at the Fermi level remains closed, and, in addition, the gap between the flat band and the remote bands closes at the $\Gamma$ point.
In the next step, we turn on the interaction with encapsulating hBN in the form of a top and bottom layer staggered sublattice potential Fig. 1(g) ($\Delta_{\rm hBN}=25$meV). This leads to the opening of the gap within the flat band. Additionally, similar to the gapped monolayer graphene case [Hunt_Ashoori_2013], a gap between the Dirac cones opens and their dispersion becomes parabolic. The remote bands peel off from the flat band at the $\Gamma$ point so that our system consists of a well defined flat band and a gap around the Fermi energy. We note that TTG with an applied electric field and hBN is, therefore, strikingly similar to the TBG/hBN structure, with an extra set of remote bands at $K$ points originating from the gapped Dirac cones. More details of these mechanisms are available in SM.
@@
We now show the Hartree-Fock quasiparticle bands for spin up and down obtained by self-consistently diagonalizing Hubbard Hamiltonian, Eq. 4. For sufficiently small $U<U_{c}$, we do not observe any strong renormalization of the bands close to the Fermi level, as shown in Fig. 2(a). The state is paramagnetic, meaning each atom has an equal spin population. On the other hand, when $U>U_{c}$, a transition to an AF ground state occurs. A clear Mott gap is opened in the flat band after the phase transition in Fig. 2(b) for $U>1.1t$. There is no gap opening for the Dirac cones, therefore, no gap is generated at the Fermi level. The AF state has the most significant spin polarization in the middle layer, which is the layer with the highest contribution to the flat band. Bilayer graphene has smaller $U_{c}$ than a monolayer, therefore the effect of $U$ is stronger in the middle of TTG. The total spin density on A and B sublattices of graphene layer building TTG is imbalanced, producing a microscopic AF state. Such AF state is analogous to an AF state in other n-layer graphene systems [Vahedi_TramblydeLaissardiere_2021, Vidarte_Lewenkopf_2024]. We note that for the system to undergo a transition from a metallic to insulating state, the presence of an hBN substrate is necessary. Without it, the system is metallic both in the paramagnetic and antiferromagnetic states, while it remains insulating for all values of U and electric field with an hBN substrate present.
The Hubbard parameter $U$ for which the para- to anti-ferro- magnetic transition occurs is lower than in the mono- and bi-layer graphene systems (Fig. 3). For example, it has been shown that in the monolayer graphene (MLG) in the Hartree-Fock approximation $U_{c}\approx 2t$ [Vahedi_TramblydeLaissardiere_2021]. We reproduced this result in our model, with small differences caused by the long-range TB hopping necessary to model TTG correctly. Equivalent calculations for bilayer graphene (BLG) yield $U^{\rm BLG}_{c}<U^{\rm MLG}_{c}$. This stems from the fact that since the dispersion in BLG is parabolic, the density of states is increased, and the ratio of interactions to kinetic energy increases as well. It is not surprising then that TBG has a smaller critical transition parameter ($U=1.1t$), as shown in Fig. 3. An extension of this Fig., showing the influence of the hBN substrate can be found in SM Fig. 7.
@@
We performed calculations for large U and concluded that the state with the lowest energy is still antiferromagnetic. Assuming ferromagnetic input (initial) state, we obtain energy difference between antiferromagnetic (ground) and ferromagnetic (higher in energy) states of $0.7193$ meV. This value is consistent in order of magnitude with TBG system, see e.g. recent works of Hou et al. [Hou_Nevidomskyy_2025] and Sánchez Sánchez et al. [Sanchez_Stauber_2025], but at odds with Adhikari et al. [Adhikari_Uchoa_2024] that claims obtaining ferromagnetic ground state at $\nu=0$ for TBG.
@@
Multiband Berry’s curvature tuning. Flat bands of twisted materials are known to realize a variety of topological phases. For example, at $\nu=-3$ filling in TBG interaction-induced Chern insulator is predicted [Sharpe_Goldhaber-Gordon_2019]. The topological phases catalog is still to be established in the TTG studied here. For a charge-neutral system for all studied electric fields, hBN strengths and Hubbard parameters $U$, flat bands near the Fermi level are always intertwined. Therefore, single band Berry’s curvature and Chern numbers are not well defined. However, multiband analogs can be defined, in our case of 2+2 bands separated from other bands by well-defined energy gaps, valid for an ample region in the parameter space of $\Delta_{\textrm{hBN}},\Delta_{V}$ and $U$. In Figs. 5 (a) and (b), we show the corresponding Hartree-Fock band structures for two representative choices of parameters (more details in SM). Details of multiband Berry’s curvature calculations are shown in the Methods section.
First, we establish that the multiband Chern number $C_{1+2+3+4}$ corresponding to all four bands is 0. This is also the case for the two flat valence (VBs) and conduction bands (CBs) separately ($C_{1+2}=C_{3+4}=0$). However, we observe a non-zero peak of mBC around $K$ and $K^{\prime}$ points, see e.g., Fig 5(c), which cancel each other out. We establish that non-zero peaks are also present around the $\Gamma$ point, in Fig. 5(c) in the corners of rhomboidal moiré mBZ. In that sense, the geometry of wavefunctions is hidden since it produces $C=0$ despite having a rich profile in the mBZ. We note that significant Berry curvature occurring at moiré K points can be intuitively understood as resulting from the Dirac cone hybridizing with the flat-bands via electric field and hBN gap. This effect provides a new scenario for the generation of topological flat-bands in trilayer moiré materials.”
@@ Discussion
Twisted trilayer graphene is a highly tunable moiré system. We predict a surprisingly small critical value of Hubbard parameter $U$ necessary to observe magnetic phase transition between para- and anti-ferromagnetic states. This AF state is microscopic, in which spins are anti-aligned on neighboring graphene A and B sites inside TTG’s AAA moiré centers. It is well-known that the AF state is not observed in pristine graphene because interactions are too weak. On the other hand, in TTG, this condition is weakened, although we note that such a state is destabilized by mirror and sublattice symmetry breaking perturbations. Effectively, we also expect stronger screening in our setup compared to that of suspended graphene, and further studies are necessary to establish optimal device design to reach the AF state in TTG. Such studies, including, e.g., quantum fluctuations, more realistic extended Hubbard models with long-range interaction, and better account for the dielectric environment and electron-electron screening models are the subject of our ongoing work. We note that the unrestricted Hartree-Fock method used in this paper has been shown sufficient to capture the ground state properties of twisted bilayer graphene [Soejima_Zaletel_2020, Potasz_MacDonald_2021, Xie_Regnault_2021, Rai_Wehling_2024], with general qualitative picture unchanged when more precise treatments of correlations, i.e. exact diagonalization, density matrix renormalization group or dynamical mean-field theory, are used. Inclusion of correlations is the subject of our future work. We note that in the absence of the hBN substrate, we did not observe the emergence of sublattice-polarized ground states. On the other hand, the detection of spontaneously broken valley symmetry [Xie_Lian_2021, Christos_Scheurer_2022], in the context of atomistic calculation, is not as straightforward, as shown in Ref. [Sanchez_Stauber_2024].
@@
Although topologically trivial at charge neutrality, TTG hosts nonzero peaks of multiband Berry’s curvature. This curvature originates from the geometry of wave functions in the flat band and can be tuned using electric fields. This can be useful in understanding low electron/hole doped phases, especially those in which carriers unequally populate valleys, i.e., valley polarized states. It would also be interesting to establish if AF and topological phases can co-exist for realistic interactions in other multilayer systems, e.g., pentalayers with hBN-induced moiré potentials, in which fractional quantum anomalous Hall states have been recently observed [Lu_Ju_2024].